OpenStudy (anonymous):

integral of sqrt (x^2+6x) dx

OpenStudy (psymon):

Try completing the square and then doing a trig substitution?

OpenStudy (anonymous):

im not sure how to do the trig substitution..

OpenStudy (anonymous):

i know the completing the square is (x+3)^2-3

zepdrix (zepdrix):

Hmm I think it's (x+3)^2-9 right? To complete the square you add and subtract 9.

zepdrix (zepdrix):

\[\large \color{royalblue}{x^2+6x+9}-9 \qquad =\qquad \color{royalblue}{(x+3)^2}-9\]

OpenStudy (psymon):

Yeah, I dont know how to do this without trig sub unfortunately, lol.

OpenStudy (anonymous):

well i dont know how to do it at all..

OpenStudy (psymon):

This calc 1 or 2?

OpenStudy (anonymous):

2

OpenStudy (psymon):

Okay, makes my life easier. So trig sub is something you should be taught then. Okay, so the idea is if you ever have a form of either: \[\sqrt{u^{2}-a^{2}}\] \[\sqrt{a^{2}-u^{2}}\]or \[\sqrt{u^{2}+a^{2}}\] where u is a variable or variable expression and a is a constant, then you can do a trig substitution. when you have the form: \[\sqrt{u^{2}-a^{2}}\], you let u = asec(theta) and the entire radical expression be atan(theta). So doing that, I can write this: \[\int\limits_{}^{}3\tan \theta du\] The next step is to solve for du. Well, we just said u was asec(theta) or just 3sec(theta), meaning du = 3sec(theta)tan(theta) giving us now: \[9\int\limits_{}^{}\tan^{2} \theta \sec \theta d \theta\]do you think you would know how to integrate that? Now keep in mind, integrating this wont give you the finished answer, but would you know how to?

OpenStudy (anonymous):

ill try. thanks

OpenStudy (psymon):

Yep. Once you integrate it, you have to go back to what our original substitutions were and use those to change the answer back into terms of x.